# Fully homomorphic encryption, outsourced databases, high-performance computing

### v1

@@ 8 Final Remark
This work proposes a new perspective on ciphertext generation in fully homomorphic encryption (FHE), shifting the focus from runtime encryption to compile-time algebraic synthesis. By precomputing a structured ciphertext basis and leveraging linearity in the encryption function, we eliminate the need for direct encryption during data ingestion, reducing overhead while preserving semantic correctness and security.
The proposed abstraction treats ciphertexts not merely as outputs of cryptographic primitives, but as composable objects synthesized from a reusable module of encrypted algebraic components. This design philosophy opens new possibilities in encrypted systems, where synthesis, slot-level layout control, and noise injection can be decoupled from cryptographic internals and expressed as part of the system architecture itself.
Beyond the specific construction introduced here, the broader message is conceptual: encryption, when viewed through the lens of algebraic programming, admits symbolic compilation strategies previously confined to plaintext computing. We believe this direction can serve as a foundation for further advances in encrypted database systems, bootstrappable encrypted pipelines, and secure compiler backends that reason about ciphertexts as algebraically structured data.
Looking ahead, an even deeper question emerges: can we eliminate encryption noise altogether, not by amortizing or refreshing it, but by constructing encryption schemes whose algebraic semantics are inherently noiseless? We suspect that such structures may lie beyond the conventional polynomial ring frameworks used in current lattice-based schemes. Candidate directions include the use of flat modules, cohomological vanishing in sheaf-theoretic encodings, or fiber product constructions that preserve exactness across encrypted morphisms. The technical barriers are considerable—extending exact sequences under encryption, controlling derived functors over non-reduced schemes, and establishing structural invariants in ciphertext categories—–but the potential theoretical payoff is equally substantial. We welcome collaborations from researchers interested in the intersection of modern algebra and cryptographic structure, and view this as a long-term invitation to reimagine encrypted computation from first mathematical principles.

### v2

@@ 8 Final Remark
This work presents a shift in the methodology of ciphertext generation within fully homomorphic encryption (FHE), transitioning from conventional runtime encryption toward a compile-time algebraic synthesis paradigm. By precomputing a structured ciphertext basis and introducing a reusable ciphertext of zero—randomized via multiplicative masking—we demonstrate that full semantic correctness and IND-CPA security can be achieved without invoking cryptographic encryption during online operation. This significantly reduces computational overhead and allows ciphertext generation to scale in parallel with plaintext throughput.
The essence of our proposal is to reframe ciphertexts as algebraically composable objects synthesized from a finite basis. In this view, the act of encryption becomes a symbolic compilation procedure: given a plaintext vector $\mathbf{m}$, the corresponding ciphertext is produced via linear synthesis over encrypted module elements and masked by scalar-randomized noise. This framing dissolves the strict separation between algebra and cryptography. Rather than treating ciphertexts as cryptographically opaque objects, our construction makes their structure explicit and manipulable by design, within the constraints of semantic security.
More broadly, our approach suggests that homomorphic ciphertexts can be understood not only as encodings of data, but as elements of a precompiled algebraic module, where synthesis, noise, and encoding layout are governed by a symbolic algebra that is deterministic at compile time and randomized only at the boundary of instantiation. This decoupling between randomness and structure allows for new architectural patterns in encrypted systems. For example, batch ciphertext ingestion pipelines can be built on top of deterministic synthesis rules, with cryptographic noise factored out as a post-hoc masking step using a single ciphertext of zero.
The implications are not merely technical. Conceptually, this work belongs to a broader school of thought that seeks to compile symbolic representations of computation directly into encrypted form, avoiding cryptographic bootstrapping at runtime and enabling algebraic reasoning over encrypted objects. We envision applications in encrypted databases, homomorphic compilers, encrypted tensor programs, and privacy-preserving stream processors, where ciphertexts are constructed by interpreters of algebraic blueprints rather than via repeated calls to cryptographic primitives.
Looking forward, we pose a natural but ambitious question: can encryption be made intrinsically noiseless, not through noise refreshing, bootstrapping, or parameter selection, but through a fundamental redesign of the algebraic semantics of the encryption function? We hypothesize that the answer may require tools from beyond traditional lattice frameworks. Potential mathematical pathways include:
  • Module-theoretic constructions over flat and faithfully exact extensions, where noise terms vanish under base change;
  • Cohomological techniques from algebraic geometry, where noise may be absorbed into acyclic sheaf resolutions or bounded derived categories;
  • Exact fibered categories over ciphertext morphisms, preserving semantic invariants under categorical encryption functors;
  • Non-commutative geometry and derived deformation theory, allowing fine-grained control over encrypted morphism noise through homotopical lifts.
These directions are speculative, but they reflect a growing realization: that ciphertexts, like programs, can be the subject of formal mathematical compilation. We believe that such investigations are not merely decorative abstractions, but necessary conceptual steps in the long-term project of building encrypted computing systems that are secure, compositional, and semantically transparent. We welcome collaboration with researchers in algebra, geometry, and category theory who seek to understand encryption not just as a cryptographic protocol, but as an algebraic transformation—one whose structure can, and should, be reasoned about with the full power of modern mathematics.